OpenStudy (anonymous):

The Richter scale is used to measure the magnitude of an earthquake. The magnitude, R, is given by the earthquake. Determine the amount of energy released be an earthquake of magnitude of 7.5. Rewrite the equation for the magnitude of an earthquake , using the magnitude of 7.5 given in the problem.

OpenStudy (tkhunny):

You'll have to provide more information. The required conversion from Richter Scale Magnitude to Energy Released is not common knowledge.

OpenStudy (anonymous):

the equation is R=0.65log(0.39E)+1.45

OpenStudy (tkhunny):

Where "E" is the energy? Okay, solve that equation for "E".

OpenStudy (anonymous):

it doesnt say to solve it, but would the new equation look like this 7.5=0.65log(0.39E) +1.45

OpenStudy (tkhunny):

Well, since you have R and not E, how else do you plan to find E? Solve it for E. Substitute 7.5 for R first, if you like, but solve for E.

OpenStudy (anonymous):

does E=-0.07

OpenStudy (tkhunny):

Not sure. I didn't work the problem. However, if we are talking about energy release, it is unlikely to be a small number and also unlikely to be negative. Solve for E and SHOW YOUR STEPS. Just one piece at a time. 7.5=0.65log(0.39E)+1.45 Subtract 1.45 7.5 - 1.45 =0.65log(0.39E) 6.05 =0.65log(0.39E) Now what?

OpenStudy (anonymous):

6.05=-0.19(0.39E) 6.05=-0.07E find LN of 6.05 1.8=-0.07E than divide both sides by -0.07 -25.71=E

OpenStudy (anonymous):

is this correct @tkhunny

OpenStudy (anonymous):

I need help solving this equation @Psysom

OpenStudy (anonymous):

This problem @Psymon its to long to write again.

OpenStudy (psymon):

So I'm guessing \[7.5 = 0.65\log(0.39E) + 1.45\] Is the starting point.

OpenStudy (anonymous):

got that but tkhunny told me to solve for E and never responded back

OpenStudy (psymon):

Well, this means we want to isolate the log term first. So I will subtract 1.45 from both sides then divide by .65.

OpenStudy (anonymous):

or subtract both sides by 1.45

OpenStudy (psymon):

Doing that gives us: \[9.31 = \log(0.39E)\] And yes, I subtracted the 1.45 first, I mentioned that :P

OpenStudy (anonymous):

so its not asking me to solve the problem it says to Rewrite the equation for the magnitude of an earthquake , using the magnitude of 7.5 given in the problem.

OpenStudy (psymon):

So now by log rules: \[x = \log_{b}M \] is the same as \[b ^{x} = M\] Now when the base is not written with the log, we assume it to be base 10. So because of our log rule, we are being told this: \[10^{9.31} = 0.39E\]

OpenStudy (anonymous):

you gave me the answers to another question i have. Simplify the problem so that it is written in the form b=loga.

OpenStudy (psymon):

Haha. Interesting. Well, what I put up there should be the answer. 10^(9.31)/.39

OpenStudy (anonymous):

always thinking a head.

OpenStudy (anonymous):

ok ya now its asking me to solve the equation.

OpenStudy (psymon):

Right, which is what I just posted :P 10^(9.31)/0.39. Pretty massive number o.o

OpenStudy (anonymous):

so i solved it wrong in the equation above.

OpenStudy (psymon):

I would say so, yeah. Somehow a -0.19 came out of nowhere O.o

OpenStudy (anonymous):

you now how it has log on the calculator i did that for 0.65.

OpenStudy (psymon):

Well, the 0.65 is just multiplying the log, it has nothing to dowith the log itself. You need to isolate the log term first. Just like you would isolate x in an algebraic equation, you want to isolatethe log term. Which means I divided both sides by 0.65.

OpenStudy (anonymous):

when i divide i get 9.38

OpenStudy (psymon):

Sorry, had to be away x_x So wait, (7.5 - 1.45)/.65 and you got 9.38?

OpenStudy (psymon):

Yeah, I still get 9.31

OpenStudy (anonymous):

ok thank u.

OpenStudy (psymon):

Mhm. And then once you have the log expression by itself, you can use the log property I mentioned to get your answer :3

OpenStudy (anonymous):

@Psymon how would u solve 10^9.31?

OpenStudy (psymon):

Calculator. Nothing much else you could do.

OpenStudy (anonymous):

i got a really big number

OpenStudy (psymon):

You should xD

OpenStudy (psymon):

And then it's divided by .35

OpenStudy (anonymous):

0.39 u mean?

OpenStudy (psymon):

Whatever the number was, sorry x_x

OpenStudy (anonymous):

its fine just in case i skiped i step or something.

OpenStudy (psymon):

Right. But no, the number is supposedtobe massive it seems. I solved for it on my calculator justto check.

OpenStudy (anonymous):

all right my final answer is 5235225499

OpenStudy (psymon):

Seemingly, yes. I mean think of it, an earthquake is a massive force, not surprised the energy level would be massive.

OpenStudy (anonymous):

got it.

OpenStudy (psymon):

cool, cool ^_^